OpenStudy (anonymous):

Shane is making ramps of different lengths. The ramps are made with rectangular and square wooden boards. The relationship between the length of the ramp in feet, f(n), and the number of square boards used, n, is represented by the function below. f(n) = 6n + 10 , where 15 ≤ n ≤ 25 Which statement is correct about the range of the length of the ramp for the given domain? It varies from 100 to 160. It varies from 100 to 150. It varies from 90 to 150. It varies from 15 to 160.

OpenStudy (anonymous):

first one

OpenStudy (across):

You're given a function: \[f(x)=6n+10\]And a domain: \[15\le n \le 25\]Just plug that domain into the function and you'll obtain: \[100 \le f(x) \le 160\]

OpenStudy (anonymous):

Thank youu soo muchh ! (: